# Compress then Serve: Serving Thousands of LoRA Adapters with Little Overhead

### v2

@@ 4 Theoretical Analysis
In this section, we seek to better understand the role of the joint diagonalization method presented in §3.1 and how this understanding further motivates the clustering approach. We will focus on the full-$\Sigma_{i}$ case with orthogonal $U$, $V$ matrices. Note that, for the same $r$, the $r\!\!-\!\!\mathsf{JD\!-\!Diag}$ has at least as large reconstruction error as $r\!\!-\!\!\mathsf{JD\!-\!Full}$ since it imposes an additional constraint on the $\Sigma_{i}$.
@@
Suppose that for all $i$, $\mathrm{rank}(B_{i}A_{i})=r_{i}$, and let
@@
Note $\max_{i}r_{i}\leq\tilde{r}\leq\sum_{i=1}^{n}r_{i}$. Then $\mathsf{JD-Full}$ (equation 2) with $r=\tilde{r}$ achieves lossless compression (perfect reconstruction), and using $r<\tilde{r}$ will give nonzero reconstruction error.
@@
Consider $n$ LoRAs ($\{A_{i},B_{i}\}_{i=1}^{n}$) with $r,n\leq d^{2}$, and form the matrix
$L=\left[\begin{array}[]{ccc}\mathrm{vec}(B_{1}A_{1})&\cdots&\mathrm{vec}(B_{n}A_{n})\end{array}\right].$
Let $\sigma_{j}$ be the singular values of $L$, sorted from largest to smallest, and let $\bar{\sigma}_{j}$ be the singular values of $\sum_{i=1}^{n}B_{i}A_{i}$. Then, using $\mathsf{JD-Full}$ (equation 2),
@@
$\frac{\sum_{i=1}^{n}\|U\Sigma_{i}V^{\top}\|^{2}_{\mathrm{Fro}}}{\sum_{i=1}^{n}\|B_{i}A_{i}\|_{\mathrm{Fro}}^{2}}\!\leq\!\frac{\sum_{j=1}^{\min(r^{2},n)}\!\!\sigma_{j}^{2}}{\sum_{j=1}^{n}\sigma_{j}^{2}}\leq 1,\>\>{and}\>\>\>\frac{\sum_{i=1}^{n}\|U\Sigma_{i}V^{\top}\!\!-\!B_{i}A_{i}\|^{2}_{\mathrm{Fro}}}{\sum_{i=1}^{n}\|B_{i}A_{i}\|_{\mathrm{Fro}}^{2}}\geq 1-\frac{\sum_{j=1}^{\min(r^{2},n)}\!\!\sigma_{j}^{2}}{\sum_{j=1}^{n}\sigma_{j}^{2}}.$
In other words, if the singular values of $L$ are not concentrated in the top $r^{2}$ entries, significant reconstruction error is unavoidable.
@@
The upper bound is smallest when the LoRAs are relatively clustered, i.e., when groups of vectors $\mathrm{vec}(B_{i}A_{i})$ are similar. This situation raises the magnitude of the largest singular values of $L$, raising the upper bound in the proposition. As the LoRAs are $d\times d$ matrices that can be thought of as points in $d^{2}$ dimensional space, for typical values of $d$ well into the hundreds, it is likely that unrelated LoRAs will be unclustered, i.e., they will have relatively low inner products with each other.
For the case of orthogonal LoRAs, the singular values of $L$ are the norms of the LoRAs, and we immediately have the following corollary: [^1] A result for isotropic Gaussian LoRAs could be obtained via the quantiles of the Marchenko-Pastur Law.
@@
Suppose (e.g., due to normalization) that the inputs to the joint diagonalization algorithm all have unit Frobenius norm, i.e., $\|B_{i}A_{i}\|_{\mathrm{Fro}}=1$. Moreover, assume that the LoRAs are all orthogonal in the sense $\mathrm{tr}((B_{i}A_{i})(B_{j}A_{j})^{\top})=0$ for $i\neq j$. Then, using the $\mathsf{JD-Full}$ method equation 2, we have $1\leq\sum_{i=1}^{n}\|\Sigma_{i}\|^{2}_{\mathrm{Fro}}\leq\min(r^{2},n),$ implying that the sum of squared Frobenius norms of the reconstructed LoRAs satisfies
@@
This implies that for the common setting where $r^{2}\ll n$, the reconstructed LoRAs will be significantly smaller than the original LoRAs and necessarily have significant reconstruction error.
The results in this section illustrate the tradeoffs of using joint diagonalization. If the LoRAs are similar or well-clustered, reconstruction error will be low. On the other hand, if the LoRAs are random and orthogonal, reconstruction error will be high.
Since the loss space of transformers is highly complex, increasing weight reconstruction error does not necessarily imply degrading LLM performance. Interestingly, in Figure 4 below, we see that while large reconstruction error rapidly decreases performance, moderate (but still relatively large, at around 0.4) reconstruction error does not damage performance and may even slightly outperform the zero-error setting. This observation motivates our focus on minimizing weight reconstruction error, while also suggesting that our approach is capable of achieving something deeper than compression. Specifically, the tendency of joint diagonalization is to find subspaces that are shared among many LoRAs when $r$ is large, and to merge subspaces when $r$ is small. When $r$ is particularly small, this tendency towards averaging all or some of the LoRA signals directly connects to the concept of merging LoRAs, whose empirical success (Shah et al., 2023; Huang et al., 2024) could explain the success of our procedure despite the nonlinearity of transformers.
Experiments in Appendix G.9 explore this idea further, comparing reconstruction of real-world LoRAs to reconstruction of randomly sampled LoRAs. The reconstruction error is generally large, but significantly lower than the reconstruction error for random noise, indicating that there is a major shared component between the LoRAs that is being successfully retained.
That said, as the number of LoRAs grows, the shared component may not be significant enough to maintain sufficiently low reconstruction error with low rank $r$. This motivates the introduction of clustering in §3.2, since clustering seeks to find groups of LoRAs that are similar and better compressible by joint diagonalization. In particular, if the number of clusters $|C_{j}|$ grows with $n$, the reconstruction error may no longer degrade with $n$ even when $r$ is fixed.
In the extreme case where $|C_{j}|=n$, each LoRA is compressed independently. By the Eckart-Young Theorem, JD applied to a single LoRA reduces to an SVD, replacing each rank-$r_{i}$ LoRA adapter $B_{i}A_{i}$ with a reduced rank-$r$ approximation, where typically $r<\frac{1}{n}\sum_{i=1}^{n}r_{i}$:
@@
As $\Sigma_{i}V_{i}^{\top}$ can be saved as a single matrix, this approach has $rn(d_{A}+d_{B})$ parameters. We refer to this $|C_{j}|=n$ method as $\mathsf{r-SVD}$ and find that it underperforms our other methods, while outperforming the baseline uncompressed LoRAs significantly. This result parallels Jiang et al. (2023b)’s observation that lowering LoRA ranks is beneficial for multi-task learning and model merging.
5 Training LoRAs & Evaluating Task Performance
@@ 5.1 Training
We trained LoRA adapters on 500 natural instruction tasks (Wang et al., 2022) using Mistral-7B-Instruct-v0.2 (Jiang et al., 2023a) as the base model. All LoRA adapters were configured with a rank of 16, i.e., $\forall i,r_{i}=16$.
We selected 10 diverse tasks (Table 2 in Appendix C) manually for consistent evaluation across experiments and randomly sampled an additional 490 tasks, resulting in a total of 500 tasks. These tasks were exclusively in English (both input and output), ensuring higher quality and thorough review (Wang et al., 2022). Each task dataset was divided into training, validation, and test sets.
Hyperparameters, such as early stopping, were tuned using the validation sets. Evaluation on the test sets demonstrated that LoRA consistently outperformed the base model in terms of both Rouge scores and loss metrics, as shown in Table 1. Details are provided in Appendix C.
Table 1: Comparison of metrics before and after LoRA training across 500 tasks.
Metric Base Model LoRA
Loss $4.99\pm 3.11$ $0.43\pm 0.57$
Exact Match $2.28\pm 7.89$ $66.66\pm 34.34$
Rouge-1 $20.38\pm 18.90$ $76.74\pm 24.89$
Rouge-L $19.66\pm 18.16$ $76.22\pm 25.27$
@@ 5.2 Evaluation
We evaluated multiple metrics for the natural instruction tasks, including cross-entropy loss, Rouge-1, Rouge-L (Lin, 2004), exact match, and agreement between uncompressed and compressed LoRA. Here, agreement measures the exact match in task-generations between the uncompressed LoRA model and the compressed LoRA model, rather than comparing to ground truth data. While detailed results and discussions for all metrics are provided in Appendix G, our primary focus in the main text is on Rouge-L. We find that all metrics correlate, but Rouge-L correlates most strongly with downstream utility. This finding aligns with prior work (Wang et al., 2022), which demonstrates that Rouge-L correlates well with classification accuracy.
@@
Joint diagonalization optimizes reconstruction error measured by the Frobenius norm, and our theoretical analysis in §4 bounds this reconstruction error. Empirically, reconstruction error and downstream Rouge-L performance correlate.
Instead of listing the absolute performance of different methods, we compute the performance difference between the base model and the LoRA model for each task. We present the ratio
@@
for the specific method in question, highlighting relative improvement with respect to the uncompressed LoRA and the base model.
@@ 6.1 Task Performance
For each method, we vary the number of $n$ LoRAs that are compressed and the compression rank $r$. We run each experiment three times with different random seeds and report the mean and standard deviation. See Table 3 for results where we evaluate on the same ten manually selected tasks (Table 2) across settings. Every compressed collection of LoRAs contains these 10 tasks (i.e., in-distribution tasks), and each collection contains the smaller collections as subsets.
@@
Figure 2 illustrates the Total Parameter Saved Ratio versus the Number of Unique LoRAs served. We only include methods that maintain over 99% of the original LoRA’s performance (as measured by RougeL). Notably, our JD methods uniquely approach the compression efficacy of a single LoRA, and with clustering, this aggressive reduction in size also maintains performance in larger LoRA collections. It is also interesting to note (see Appendix G) that JD variants often increase generalization and outperform the original LoRA.
For efficiency, we limited the JD methods to ten iterations instead of pursuing full convergence. While the alternating algorithm quickly reaches an approximation of the minimizer, squeezing out the last few digits of precision takes many more iterations with limited to no performance gain. Appendix G.10 also evaluates an alternative eigenvalue iteration algorithm that more rapidly converges once $U,V$ are close to a minimizer, with minimal performance differences.
Figure 2: Total Parameter Saved Ratio versus the Number of Unique LoRAs served. We only include methods that maintain over 99% of the original LoRA’s performance (as measured by RougeL). The Total Parameter Saved Ratio depicts the number of parameters saved for a system with a large number $n$ of different LoRAs. It is computed as: $r_{total}:=1-\frac{\text{num. parameters after compression}}{\text{num. parameters before compression}}$.
@@ 6.2 Performance and Reconstruction Error
Figure 4 relates reconstruction error and performance. The $y$-axis measures the mean performance improvement of Rouge-L relative to uncompressed LoRA, and the $x$-axis quantifies the mean reconstruction error between the compressed reconstruction of the product $BA$ and the original uncompressed product $BA$. Although the relationship between performance and reconstruction error is nonlinear, it demonstrates a generally decreasing, somewhat exponential trend. Notably, the minimal reconstruction error does not correlate with optimal performance, indicating that a degree of lossy reconstruction may be advantageous for enhancing generalization.
To select hyperparameters (compression rank and number of clusters) for the clustering experiments, we first assessed reconstruction error on a single LoRA module over a range of settings (see Appendix F). These preliminary experiments enabled efficient selection of cluster counts and rank values for compressing all LoRA modules.
@@
Figure 4: Throughput ratio when serving varying numbers of LoRAs with vLLM.
@@ 6.3 Throughput of Serving Compressed LoRAs
Results in the previous sections demonstrate how to select an appropriate joint compression setting guided by the reconstruction error, such that the performance of the original LoRAs is preserved. Naturally, the rank and/or the number of clusters for the compression needs to increase as we compress larger LoRA collections to match LoRA performance.
@@
In Figure 4 we study how throughput with various compression settings compares to the vLLM multi-LoRA throughput with the matched GPU memory footprint. Specifically, for each number of unique LoRAs served and each compression setting, we compute the corresponding number of LoRAs to be placed on the GPU during serving and report the ratio of the two throughputs. For example, when serving 64 unique LoRAs and using rank 64 JD-Full compression, we report the ratio of throughputs of rank 64 JD-Full and vLLM multi-LoRA with 6 LoRAs allowed on the GPU at a time (see Appendix E for details). As the number of unique LoRAs increases, vLLM multi-LoRA throughput degrades as it needs to schedule the requests and load and offload the adapters. We note that vLLM multi-LoRA already employs many advanced system optimization techniques, such as efficient scheduling and non-blocking CPU-GPU communication when swapping LoRAs (Sheng et al., 2023; Kwon et al., 2023), but system optimization alone is not sufficient to fully mitigate throughput degradation when serving many LoRAs.
@@
In Figure 4 we see that across all LoRA collection sizes our compression techniques improve the throughput of vLLM multi-LoRA. Additionally, we highlight regions for each compression setting where compression is sufficiently moderate to achieve 99%+ of LoRA performance, according to the results in Section 6.2. We also note that compression with a larger rank or too many clusters does not improve baseline throughput when serving a smaller number of LoRAs and should not be used in such cases. For example, rank 16 JD-Full improves baseline throughput with 4 and 8 LoRAs, but will underperform with more LoRAs, while 7 cluster rank 64 JD-Full does not improve throughput with 64 or fewer LoRAs, but when serving 1000+ LoRAs it improves the throughput significantly while maintaining the performance. To conclude, an appropriate joint compression setting improves vLLM multi-LoRA throughput and preserves performance for LoRA collections of any size between 4 and 1024, as we showed in Figure 1. Specific compression settings for each LoRA collection size are listed in Appendix E.
@@
Finally, we note that vLLM extensively uses custom CUDA kernels. To accommodate our compression techniques, we minimally adjusted the vLLM code to generate additional kernels needed by the compressed LoRAs. There likely is room for improvement to optimize the newly added kernels.
Additional details
@@
In this experiment we considered a varying number of rank-16 LoRAs, using a dataset of Shakespeare sonnets as inputs [^2] https://www.kaggle.com/datasets/shivamshinde123/william-shakespeares-sonnet/data arriving asynchronously. We measured throughput, i.e., the number of requests served per second when generating ten tokens per request. The base model was Mistral 7B Instruct as in the other experiments; we simulated random LoRAs and assigned inputs to LoRAs at random. Experiments were conducted on H100 80GB GPU capped at 40% memory consumption. This was done to reflect cost concerns in practical situations where a service provider might want to serve many LoRAs from cheaper hardware with lower memory than higher-end GPUs. This setting also takes into account the scenario where the LLM is large compared to the size of GPU and yet a provider may want to serve many LoRAs efficiently using the same device.

### v3

@@ 4 Theoretical Analysis
In this section, we seek to better understand the role of the joint diagonalization method in §3.1 and how it motivates the clustering approach. We focus on the full-$\Sigma_{i}$ case with orthogonal $U$, $V$ matrices. Note that, for the same $r$, the $r$-JD-Diag has at least as large reconstruction error as $r$-JD-Full since it imposes an additional constraint on the $\Sigma_{i}$.
@@
Suppose $\mathrm{rank}(B_{i}A_{i})=r_{i}$ for all $i$, and let
@@
Note $\max_{i}r_{i}\leq\tilde{r}\leq\sum_{i=1}^{n}r_{i}$. Then JD-Full (equation 2) with $r=\tilde{r}$ compresses losslessly (perfect reconstruction), while $r<\tilde{r}$ will give nonzero reconstruction error.
@@
Consider $n$ LoRAs $\{A_{i},B_{i}\}_{i=1}^{n}$ with $r,n\leq d^{2}$, and form the matrix $L=\left[\begin{array}[]{ccc}\mathrm{vec}(B_{1}A_{1})&\cdots&\mathrm{vec}(B_{n}A_{n})\end{array}\right].$ Let $\sigma_{j}$ be the singular values of $L$, sorted from largest to smallest, and let $\bar{\sigma}_{j}$ be the singular values of $\sum_{i=1}^{n}B_{i}A_{i}$. Then, using JD-Full (equation 2),
@@
$\frac{\sum_{i=1}^{n}\|U\Sigma_{i}V^{\top}\|^{2}_{\mathrm{Fro}}}{\sum_{i=1}^{n}\|B_{i}A_{i}\|_{\mathrm{Fro}}^{2}}\!\leq\!\frac{\sum_{j=1}^{\min(r^{2},n)}\!\!\sigma_{j}^{2}}{\sum_{j=1}^{n}\sigma_{j}^{2}}\leq 1,\textrm{ and }$
$\frac{\sum_{i=1}^{n}\|U\Sigma_{i}V^{\top}\!\!-\!B_{i}A_{i}\|^{2}_{\mathrm{Fro}}}{\sum_{i=1}^{n}\|B_{i}A_{i}\|_{\mathrm{Fro}}^{2}}\geq 1-\frac{\sum_{j=1}^{\min(r^{2},n)}\!\!\sigma_{j}^{2}}{\sum_{j=1}^{n}\sigma_{j}^{2}}.$
In other words, reconstruction error is unavoidable if $L$’s singular values are not concentrated in the top $r^{2}$ entries.
@@
The upper bound is smallest when the LoRAs are relatively clustered, i.e., when groups of vectors $\mathrm{vec}(B_{i}A_{i})$ are similar. This situation raises the magnitude of the largest singular values of $L$, raising the upper bound in the proposition. As the LoRAs are $d\times d$ matrices that can be thought of as points in $\mathbb{R}^{d^{2}}$, for typical values of $d$ well into the hundreds, it is likely that unrelated LoRAs will be unclustered, i.e., they will have relatively low inner products with each other.
For orthogonal LoRAs, the singular values of $L$ are the norms of the LoRAs, suggesting the following corollary: [^2] A result for isotropic Gaussian LoRAs could be obtained via the quantiles of the Marchenko-Pastur Law.
@@
Suppose (e.g., due to normalization) that the inputs to the joint diagonalization algorithm all have unit Frobenius norm, i.e., $\|B_{i}A_{i}\|_{\mathrm{Fro}}=1$. Moreover, assume that the LoRAs are all orthogonal in the sense $\mathrm{tr}((B_{i}A_{i})(B_{j}A_{j})^{\top})=0$ for $i\neq j$. Then, using the JD-Full method equation 2, we have $1\leq\sum_{i=1}^{n}\|\Sigma_{i}\|^{2}_{\mathrm{Fro}}\leq\min(r^{2},n),$ implying that the sum of squared Frobenius norms of the reconstructed LoRAs satisfies
@@
This implies that for the common setting where $r^{2}\ll n$, the reconstructed LoRAs will be significantly smaller than the original LoRAs, with significant reconstruction error.
Our analysis illustrates the tradeoffs of joint diagonalization. If the LoRAs are similar or well-clustered, reconstruction error will be low. On the other hand, if the LoRAs are random and orthogonal, reconstruction error will be high.
Since the loss space of transformers is highly complex, increasing reconstruction error does not necessarily degrade LLM performance. Interestingly, Figure 3 below shows that while large reconstruction error rapidly decreases performance, moderate (but still relatively large, at around 60%) reconstruction error does not damage performance and may even slightly outperform the zero-error setting. At the same reconstruction error, clustering outperforms non-clustering. This motivates our focus on minimizing reconstruction error, while also suggesting that our approach achieves something deeper than compression. Specifically, joint diagonalization finds subspaces that are shared among many LoRAs when $r$ is large and merges subspaces when $r$ is small. When $r$ is particularly small, this tendency towards averaging all or some of the LoRAs connects to merging LoRAs, whose empirical success (Shah et al., 2023; Huang et al., 2024) could explain the procedure’s success despite the nonlinearity of transformers.
Appendix H.11 explores this idea further, comparing reconstruction of real-world LoRAs to reconstruction of randomly sampled LoRAs. The reconstruction error is generally large, but significantly lower than the reconstruction error for random noise, indicating that a major shared component between the LoRAs is successfully retained.
That said, as the number of LoRAs grows, the shared component may not be significant enough to maintain sufficiently low reconstruction error with low rank $r$. This motivates the introduction of clustering in §3.2, since clustering seeks to find groups of LoRAs that are similar and better compressible by joint diagonalization. In particular, if the number of clusters $k$ grows with $n$, the reconstruction error may no longer degrade with $n$ even when $r$ is fixed.
In the extreme case where $k=n$, each LoRA is compressed independently. By the Eckart-Young Theorem, JD applied to a single LoRA reduces to an SVD, replacing each rank-$r_{i}$ LoRA adapter $B_{i}A_{i}$ with a reduced rank-$r$ approximation, where typically $r<\frac{1}{n}\sum_{i=1}^{n}r_{i}$:
@@
As $\Sigma_{i}V_{i}^{\top}$ can be saved as a single matrix, this approach has $rn(d_{A}+d_{B})$ parameters. We refer to this $k=n$ method as r-SVD and find that it underperforms our other methods while slightly outperforming the baseline uncompressed LoRAs. This result parallels Jiang et al. (2023b)’s observation that lowering LoRA ranks is beneficial for multi-task learning and model merging.
5 Training & Performance Evaluation
@@ 5.1 Training
We trained LoRA adapters on 1000 natural instruction tasks (Wang et al., 2022) using Mistral-7B-Instruct-v0.2 (Jiang et al., 2023a) as the base. We set all LoRA adapter ranks to 16 (i.e., $\forall i,r_{i}=16$), except for those in our ablation study (Appendix H.1), where we vary the LoRA rank.
We selected 10 diverse tasks (Table 2 in Appendix C) manually for consistent evaluation across experiments and randomly sampled an additional 990 tasks, resulting in a total of 1000 tasks (Table 3). The tasks went through a robust reviewing protocol to ensure high quality and diversity. Each task data was divided into training, validation, and test sets.
Hyperparameters, such as early stopping, were tuned using the validation sets. Table 1, Appendix C shows that on the test sets, LoRA consistently outperformed the base model in terms of Rouge scores and loss metrics.
@@ 5.2 Evaluation
We evaluated multiple metrics for the natural instruction tasks, including cross-entropy loss, Rouge-1, Rouge-L (Lin, 2004), exact match, and agreement between uncompressed and compressed LoRA. Here, agreement measures the exact match in task-generations between the uncompressed LoRA model and the compressed LoRA model, rather than comparing to ground truth data. While detailed results and discussions for all metrics are provided in Appendix H, our primary focus in the main text is on Rouge-L. We find that all metrics correlate, but Rouge-L correlates most strongly with downstream utility. This finding aligns with prior work (Wang et al., 2022), which demonstrates that Rouge-L correlates well with classification accuracy.
@@
Joint diagonalization optimizes reconstruction error measured by the Frobenius norm, bounded by our theoretical analysis in §4. We empirically study the relation between the reconstruction error and downstream Rouge-L performance in Section 6.2.
Instead of listing absolute performance, we compute the performance difference between the base model and the LoRA model for each task via the ratio
@@
for the method in question, highlighting relative improvement wrt the uncompressed LoRAs.
@@ 6.1 Task Performance
For each method, we vary the number $n$ of compressed LoRAs and the compression rank $r$. We run each experiment three times with different random seeds and report the mean and standard deviation. See Table 7 for results evaluated on the same ten manually-selected tasks (Table 2) across settings. Every compressed collection of LoRAs contains these 10 tasks (i.e., in-distribution tasks), and each collection contains the smaller collections as subsets.
@@
Figure 2 illustrates the Rouge-L scores of the compressed LoRAs divided by the Rouge-L scores of the uncompressed LoRAs. JD variants often increase generalization and outperform the original LoRA. Notably, our JD methods approach the compression efficacy of a single LoRA, and with clustering, this aggressive reduction in size also maintains performance in larger collections. Appendix H includes tables of additional relative and absolute metrics.
For efficiency, we limited the JD methods to ten iterations instead of full convergence. While the alternating algorithm quickly reaches an approximate minimizer, squeezing out the last few digits of precision takes many more iterations with limited to no performance gain. Appendix H.12 also evaluates an alternative iterative algorithm that converges more rapidly once $U,V$ are close to a minimizer, with minimal performance differences.
Figure 2: Performance after compression. We compare the performance of compressed LoRAs relative to uncompressed ones, with higher values on both axes reflecting better performance. The Total Parameter Saved Ratio depicts the number of parameters saved for a system with a large number $n$ of different LoRAs. It is computed as: $r_{total}:=1-\frac{\text{num. parameters after compression}}{\text{num. parameters before compression}}$.
@@ 6.2 Performance and Reconstruction Error
Figure 3 relates reconstruction error and performance. The $y$-axis measures the mean performance improvement of Rouge-L relative to uncompressed LoRA, and the $x$-axis quantifies the mean relative reconstruction error between the compressed reconstruction of the product $BA$ and the original product $BA$. Although performance and reconstruction error relate non-linearly, we see a decreasing, somewhat exponential trend. Notably, minimizing reconstruction error does not yield optimal performance, indicating that mild lossy reconstruction may enhance generalization. Interestingly, under the clustering approach, compared to non-clustering, even more aggressive lossy reconstruction can outperform less lossy reconstruction, suggesting that reconstruction error is even less critical for performance in the clustering scenario.
@@
To select hyperparameters (compression rank and number of clusters) for the clustering experiments, we first assessed reconstruction error on a single LoRA module over a range of settings (see Appendix G). These preliminary experiments enabled efficient selection of cluster counts and rank values for compressing all LoRA modules.
6.3 Benefits of Compression
Compressing LoRAs reduces their parameter counts, thus lowering their overall memory footprint. While this offers many benefits, in both training and inference scenarios parameters are often transferred between different memory hierarchies (e.g., from CPU to GPU), and these transfers usually scale linearly with the amount of data moved. At the same time, compressed LoRAs alter the forward-pass formulation (unless merged with the base model weights). As shown in Figure 5 in the Appendix E, although compression greatly reduces memory usage and transfer time, it does not affect the forward-pass latency. Of the various ways to leverage these improvements, this work focuses on optimizing inference for multiple LoRAs using vLLM.
@@ 6.4 Throughput of Serving Compressed LoRAs
The previous sections demonstrate how to select an appropriate joint compression setting guided by the reconstruction error, such that the performance of the original LoRAs is preserved. Naturally, the rank and/or the number of clusters for the compression needs to increase as we compress larger LoRA collections to match LoRA performance.
@@
Figure 4 studies how throughput with various compression settings compares to the vLLM multi-LoRA throughput with the matched GPU memory footprint. Specifically, for each number of unique LoRAs served and each compression setting, we compute the corresponding number of LoRAs to be placed on the GPU during serving and report the ratio of the two throughputs. For example, when serving 64 unique LoRAs and using rank 64 JD-Full compression, we report the ratio of throughputs of rank 64 JD-Full and vLLM multi-LoRA with 6 LoRAs allowed on the GPU at a time (see Appendix F for details). As the number of unique LoRAs increases, vLLM multi-LoRA throughput degrades as it needs to schedule the requests and load and offload the adapters. We note that vLLM multi-LoRA already employs advanced optimizations, such as efficient scheduling and non-blocking CPU-GPU communication when swapping LoRAs as well as techniques introduced in S-LoRA (Sheng et al., 2023; Kwon et al., 2023), but system optimization alone is insufficient to mitigate throughput degradation when serving many LoRAs.
@@
Figure 4 shows that across LoRA collection sizes our compression techniques improve the throughput of vLLM multi-LoRA. Additionally, we highlight regions for each compression setting where compression is sufficiently moderate to achieve 99%+ of LoRA performance, according to the results in §6.2. Compression with a larger rank or too many clusters does not improve baseline throughput when serving a smaller number of LoRAs and should not be used in such cases. For example, rank 16 JD-Full improves baseline throughput with 4 and 8 LoRAs, but will underperform with more LoRAs, while 25 clusters rank 15 JD-Full does not improve throughput with 32 or fewer LoRAs, but when serving 1000+ LoRAs it improves the throughput significantly while maintaining the performance. Overall, an appropriate joint compression setting improves vLLM multi-LoRA throughput and preserves performance for LoRA collections of any size between 4 and 1024, as in Figure 1. Appendix F provides compression settings for each collection size.
@@
vLLM extensively uses custom CUDA kernels. To accommodate our compression techniques, we minimally adjusted the vLLM code to generate additional kernels needed by the compressed LoRAs and used the Punica (Chen et al., 2023) kernel to further accelerate matrix multiplication. Pseudocode is given in §F.4 to show how we use the batch multiplication kernel. There likely is room for improvement to optimize the newly added kernels.
Figure 4: Throughput ratio when serving varying numbers of LoRAs with vLLM. Highlighted settings preserve at least 99% of the uncompressed LoRA performance.
Additional details.
@@
In this experiment, we considered a varying number of rank-16 LoRAs, using a dataset of Shakespeare sonnets as inputs [^3] https://www.kaggle.com/datasets/shivamshinde123/william-shakespeares-sonnet/data arriving asynchronously. We measured throughput, i.e., the number of requests served per second when generating ten tokens per request. The base was Mistral 7B Instruct; we simulated random LoRAs and assigned inputs to LoRAs at random. Experiments were conducted on H100 80GB GPU capped at 40% memory consumption to reflect situations where a service provider might want to serve many LoRAs from cheaper hardware with lower memory than higher-end GPUs. This setting applies to the scenario where the LLM is large compared to the size of GPU and yet a provider may want to serve many LoRAs efficiently using one device.
6.5 Recommendations
JD-Full is generally preferred over JD-Diag, although for smaller numbers of LoRAs (less than 100), the performance difference is negligible. While JD-Full alone is effective up to 100 LoRAs, incorporating clustering at scales of 500-1000 LoRAs significantly enhances performance.
We recommend the following procedure for hyperparameter selection. For $\leq 100$ LoRAs, JD-Full can be used without substantial degradation, using a rank $\approx(\text{number of LoRAs}/2)+7$. Beyond 100 LoRAs, clustering becomes increasingly critical. A robust method for any number of LoRAs up to 1000 uses JD-Full with clustering. Specifically, select a LoRA module from the middle of the network, apply a compression rank of 16, and experiment with an exponentially increasing number of clusters. Compute the reconstruction error for each setting on this module across all LoRAs—a computationally efficient process. Choose the minimal number of clusters that achieves a reconstruction loss below 0.6, and then use these settings across LoRA modules. Figure 6 in the Appendix illustrates this procedure applied to 500 LoRAs.
Tuning hyperparameters as discussed above using reconstruction loss as a validation metric is convenient since it can be done efficiently on CPU without expensive LLM evaluation. As our experiments demonstrate, compression settings that achieve below 0.6 reconstruction loss reliably preserve 99% or more of the LoRA performance, sometimes even outperforming the original LoRAs.
For inference, this procedure is executed as a preprocessing step before deploying our inference server. As new LoRAs are submitted, they are initially served uncompressed. A background CPU job can periodically re-run the compression algorithm and update the served LoRA parameters with the compressed versions.